# Looking for the $\gamma$-Ray Cascades of the KM3-230213A Neutrino Source

### v1

@@ 2.2 Fermi-LAT $\gamma$-ray Analysis
We define the region-of-interest (RoI, which is defined differently from that of KM3NeT, as the term is typically used to describe the region used to fit the $\gamma$-ray background) for our $\gamma$-ray study as a $15^{\circ}$ radius centered on the reported best-fit neutrino position from KM3NeT (RA = 94.3, Dec = $-7.8^{\circ}$). To account for localization uncertainties in the neutrino position, we perform a systematic grid search spanning ±$\pm 2^{\circ}$ around the KM3NeT coordinates in increments of $0.2^{\circ}$. This choice of spatial increment is informed by the typical correlation scale of Test Statistic (TS) values for faint LAT point sources [25]. Here, the TS quantifies the source detection significance and is defined as $\text{TS}=-2\log(\mathcal{L}{{}_{\text{null}}}/\mathcal{L}_{\text{alt}})$, where $\mathcal{L}_{\text{alt}}$ and $\mathcal{L}_{\text{null}}$ denote the likelihood values with and without the source, respectively. In this context, the TS value serves as a measure of the statistical significance of a source, and is approximately related to the detection significance in units of standard deviation, $\sigma$, by $\sqrt{\text{TS}}\approx\sigma$ [26]. The analysis is conducted consistently across various temporal intervals—weekly, monthly, yearly, and over the full dataset—to investigate potential steady and transient emission scenarios.
@@
First, we optimize and fit the RoI likelihood model. After the initial fit, we further optimize the model by freeing the normalization parameters of the diffuse background components as well as the parameters of nearby sources. Specifically, we free the normalization parameter for all sources within 5 of the RoI center. To search for additional unknown sources, we run the fermipy source-finder algorithm, iteratively identifying potential new point sources in the RoI with a significance threshold of $\sqrt{\text{TS}}>5$.
The spectral energy distribution (SED) for candidate sources is derived by independently fitting fluxes in each energy bin, assuming a power-law spectrum with an index $\Gamma=0.9$ (a choice justified by the hardness of the cascade spectrum shown in Fig. 1 for a time delay of 1 year). We note, however, that choosing different $\gamma$-ray spectra “inside" each individual energy bin has a very small effect on our analysis, due to the small width of each bin. For energy bins lacking significant excess emission, we compute flux upper limits at a 95% confidence level (CL) using the delta-log-likelihood method. An example of an SED curve upper limit assuming a power-law index of $\Gamma=0.9$ is shown in Fig. 2.
@@ 3 Results
We perform a comprehensive, spatially and temporally resolved search for electromagnetic counterparts to the ultra-high-energy neutrino event KM3-230213A utilizing nearly 17 years of Fermi-LAT data to constrain associated $\gamma$-ray emission predicted by the hadronic cascade model shown in Fig. 1. Our analysis involves scanning a region within ±$\pm 2^{\circ}$ of the best-fit neutrino event position in increments of 0.2 and examining multiple time scales (weekly, monthly, yearly, and cumulative intervals) from August 2008 to February 2025. We report no statistically significant $\gamma$-ray detection at the position of KM3-230213A during any of the analyzed time intervals (Figs. 3 and 4).
@@
Figure 5: Constraints on the IGMF strength as a function of the redshift to the putative KM3NeT source, derived assuming no detection of the cascaded $\gamma$-ray flux within periods of 1 year (black), 2 years (cyan), or 7 years (green-dashed) after the neutrino event. Shaded regions indicate parameter space excluded by our analysis. Constraints are shown at the best-fit location of the KM3NeT event, but change only negligibly over the angular uncertainty region of the event. The 1- and 2-year constraints rely on the small-angle approximation ($\delta\ll\pi$), assuming minimal $\gamma$-ray deflection. We demonstrate that extending the observation window to 7 years following the KM3-230213A (i.e., 5 more years of Fermi-LAT observations) reaches the limit of this approximation ($\sim 10^{-13}$ G).
@@
In addition, our analysis demonstrates that extending the Fermi-LAT observation period by approximately 5 more years pushes us to the limit of the small-angle approximation regime used in our current cascade modeling. This extension necessitates the development of more sophisticated theoretical models to accurately describe the increased angular dispersion of cascade photons. As such, continued Fermi-LAT observations over the upcoming years will provide critical insight into the cascade behavior beyond the small-angle approximation, thereby allowing us to more robustly characterize the IGMF properties.
@@
These results indicate that the emission from 4FGL J0616.2-0653 is most consistent with steady diffuse emission or other soft-spectra sources (e.g., star-forming galaxies), rather than transient or steady cascades associated with the KM3NeT neutrino event.
@@ 3.2 A New Fermi-LAT Source near the KM3NeT Event
The analysis of the full Fermi-LAT dataset reveal a $\gamma$-ray excess at a location previously unassociated with any known $\gamma$-ray source [27]. The source is significantly brighter during the period between 2010–2011 than during the remainder of the analysis period. Among the candidates listed in [3], the source NVSS J060639–063421 is the closest to the identified $\gamma$-ray excess. Adopting the 17-year best-fit localization at (RA $=91.97^{\circ}$, Dec $=-6.47^{\circ}$, R(68%) = 0.4) or the 2010–11 localization at (RA $=92.14^{\circ}$, Dec $=-6.42^{\circ}$, R(68%)=), NVSS J060639–063421, with coordinates (RA $\sim 91.67^{\circ}$, Dec $\sim-6.57^{\circ}$), lies at angular distances of approximately $0.3^{\circ}$ and $0.5^{\circ}$, respectively, making it the most plausible candidate counterpart to the $\gamma$-ray excess.
@@
Figure 7: Analogous to Fig. 6, but for the newly identified Fermi-LAT source located $\sim 2^{\circ}$ from the best-fit position of KM3-230213A. We highlight a significant spectral hardening observed specifically during 2010–2011, potentially indicative of an earlier transient event consistent with low intergalactic magnetic field scenarios.
To further investigate the nature of this source, we perform a time-resolved spectral analysis, splitting the dataset into yearly intervals. We find that, while the source is typically faint, it exhibits a significant spectral hardening during the 2010–2011 time frame, with a best-fit power-law index of $\Gamma\sim 1.8$, compared to its long-term average of $\Gamma\sim 2.3$ as shown in Fig. 7. This spectral change may suggest a possible transient episode of high-energy activity, which could be associated with a past acceleration event. Further constraints on the source’s variability, especially in multiwavelength data, is necessary to pin down the nature of this emission.
@@ 4.1 Assumptions in Modeling the Cascade
The simplifying model we adopt here to compute the $\gamma$-ray spectra is limited by a few assumptions. While the cascades become co-linear with the primary emission in the limit of vanishing magnetic fields, large magnetic fields can push the cascades out to several tens of degrees. In modeling the electromagnetic cascades, we assume that the angular extension of the cascade is smaller than $\sim 1^{\circ}$, which is true for time delays that are small enough to observe on $\mathcal{O}(1-10)$ yr time scales. We also assume a simple redshift dependence of the magnetic field, which becomes important for redshifts greater than about $z\sim$1. However, our strongest results are found at smaller redshifts, where this dependence has vanishing effects. Finally, we adopt the EBL model of Saldana-Lopez, et al. [22]. We find that using either the ±$\pm 1\sigma$ fits to the EBL from this model does not significantly change our results.
@@ 4.2 The True Ultra-High Energy Neutrino Flux
Throughout this paper, we have assumed that the KM3NeT event was produced by a putative source that not only produced one event ($k$=1), but also has a neutrino flux which was expected to produce one event ($\lambda$=1). This approach matches the recent work of Ref. [23], who utilized an identical approach to calculate the expected GeV and TeV $\gamma$-ray signals.
@@ 5 Conclusions
One potential caveat to the above constraint stems from the possibility that the KM3NeT source is a “hidden" neutrino source, where the electromagnetic radiation from the central engine is efficiently absorbed by the dense surrounding medium [18]. In such scenarios, the high-energy $\gamma$-ray emission is effectively absorbed and re-emitted at lower and lower energies, eventually falling below the Fermi-LAT $\gamma$-ray band. However, such sources require extremely high x-ray luminosities in order to effectively absorb the high-energy emission, making them highly detectable by current x-ray instrumentation. Moreover, the cascading high-energy emission must eventually escape the hidden source, typically at MeV-energies. The detection of high energy neutrino sources without detectable GeV $\gamma$-ray counterparts thus provides a strong motivation supporting future MeV $\gamma$-ray instruments such as AMEGO [34] and e-ASTROGAM [35].

### v2

@@ 2.2 Fermi-LAT $\gamma$-ray Analysis
We define the region-of-interest (RoI, which is defined differently from that of KM3NeT, as the term is typically used to describe the region used to fit the $\gamma$-ray background) for our $\gamma$-ray study as a $15^{\circ}$ radius centered on the reported best-fit neutrino position from KM3NeT (RA = 94.3, Dec = $7.8^{\circ}$). To account for localization uncertainties in the neutrino position, we perform a systematic grid search spanning $\pm 2^{\circ}$ around the KM3NeT coordinates in increments of $0.2^{\circ}$. This choice of spatial increment is informed by the typical correlation scale of Test Statistic (TS) values for faint LAT point sources [36]. Here, the TS quantifies the source detection significance and is defined as $\text{TS}=-2\log(\mathcal{L}{{}_{\text{null}}}/\mathcal{L}_{\text{alt}})$, where $\mathcal{L}_{\text{alt}}$ and $\mathcal{L}_{\text{null}}$ denote the likelihood values with and without the source, respectively. In this context, the TS value serves as a measure of the statistical significance of a source, and is approximately related to the detection significance in units of standard deviation, $\sigma$, by $\sqrt{\text{TS}}\approx\sigma$ [37]. The analysis is conducted consistently across various temporal intervals—weekly, monthly, yearly, and over the full dataset—to investigate potential steady and transient emission scenarios.
@@
First, we optimize and fit the RoI likelihood model. After the initial fit, we further optimize the model by freeing the normalization parameters of the diffuse background components as well as the parameters of nearby sources. Specifically, we free the normalization parameter for all sources within 5 of the RoI center.
The spectral energy distribution (SED) for candidate sources is derived by independently fitting fluxes in each energy bin, assuming a power-law spectrum with an index $\Gamma=0.9$ (a choice justified by the hardness of the cascade spectrum shown in Fig. 1 for a time delay of 1 year). We note, however, that choosing different $\gamma$-ray spectra “inside” each individual energy bin has a very small effect on our analysis, due to the small width of each bin. For energy bins lacking significant excess emission, we compute flux upper limits at a 95% confidence level (CL) using the delta-log-likelihood method. An example of an SED curve upper limit assuming a power-law index of $\Gamma=0.9$ is shown in Fig. 2.
@@
Figure 5: Constraints on the IGMF strength as a function of the redshift to the putative KM3NeT source, derived assuming IGMF evolution with $\alpha=4$ and no detection of the cascaded $\gamma$-ray flux within periods of 1 year (black), 2 years (cyan), or 7 years (green-dashed) after the neutrino event. Shaded regions indicate parameter space excluded by our analysis. Constraints are shown at the best-fit location of the KM3NeT event, but change only negligibly over the angular uncertainty region of the event. The 1- and 2-year constraints rely on the small-angle approximation ($\delta\ll\pi$), assuming minimal $\gamma$-ray deflection. We demonstrate that extending the observation window to 7 years following the KM3-230213A (i.e., 5 more years of Fermi-LAT observations) reaches the limit of this approximation ($\sim 10^{-13}$ G).
@@ 3 Results
We perform a comprehensive, spatially and temporally resolved search for electromagnetic counterparts to the ultra-high-energy neutrino event KM3-230213A utilizing nearly 17 years of Fermi-LAT data to constrain associated $\gamma$-ray emission predicted by the hadronic cascade model shown in Fig. 1. Our analysis involves scanning a region within $\pm 2^{\circ}$ of the best-fit neutrino event position in increments of 0.2 and examining multiple time scales (weekly, monthly, yearly, and cumulative intervals) from August 2008 to February 2025. We report no statistically significant $\gamma$-ray detection at the position of KM3-230213A during any of the analyzed time intervals (Figs. 3 and 4).
@@
In addition, our analysis demonstrates that extending the Fermi‑LAT observation period by approximately 5 more years pushes us to the limit of the small-angle approximation regime used in our current cascade modeling. This extension necessitates the development of more sophisticated theoretical models to accurately describe the increased angular dispersion of cascade photons. As such, continued Fermi‑LAT observations over the upcoming years will provide critical insight into the cascade behavior beyond the small-angle approximation, thereby allowing us to more robustly characterize the IGMF properties.
@@
These results indicate that the emission from 4FGL J0616.2-0653 is most consistent with steady diffuse emission or other soft-spectra sources (e.g., star-forming galaxies). While the persistently soft spectrum disfavors electromagnetic cascades as the dominant emission mechanism, to definitively exclude 4FGL J0616.2-0653 as the KM3NeT neutrino source would require a redshift measurement in order to determine whether underdeveloped cascade scenario could be relevant.
@@ 3.2 A New Fermi-LAT Source near the KM3NeT Event
The analysis of the full Fermi-LAT dataset reveals a $\gamma$-ray excess at a location previously unassociated with any known $\gamma$-ray source [38]. The source is significantly brighter during the period between 2010–2011 than during the remainder of the analysis period. Among the candidates listed in [3], the source NVSS J060639–063421 is the closest to the identified $\gamma$-ray excess. Adopting the 17‐year best-fit localization at (RA $91.97^{\circ}$, Dec $6.47^{\circ}$, R(68%) = 0.4) or the 2010–11 localization at (RA $92.14^{\circ}$, Dec $6.42^{\circ}$, R(68%)= 0.5), NVSS J060639–063421, with coordinates (RA $\sim 91.67^{\circ}$, Dec $\sim-6.57^{\circ}$), lies at angular distances of approximately $0.3^{\circ}$ and $0.5^{\circ}$, respectively, making it the most plausible candidate counterpart to the $\gamma$-ray excess.
@@
Figure 7: Analogous to Fig. 6, but for the newly identified Fermi-LAT source located $\sim 2^{\circ}$ from the best-fit position of KM3-230213A. We highlight a significant spectral hardening observed specifically during 2010–2011 (top right), potentially indicative of an earlier transient event consistent with low intergalactic magnetic field scenarios. The TS for the emission integrated over the year following the KM3NeT neutrino event (2023–24) is shown in the bottom row, where the spectral index reaches the softest allowed value.
To further investigate the nature of this source, we perform a time-resolved spectral analysis, splitting the dataset into yearly intervals. We find that, while the source is typically faint, it exhibits a significant spectral hardening during the 2010–2011 time frame, with a best-fit power-law index of $\Gamma\sim 1.8$, compared to its long-term average of $\Gamma\sim 2.3$ as shown in Fig. 7. This spectral change may suggest a possible transient episode of high-energy activity, which could be associated with a past acceleration event. While the 2010–11 spectral hardening occurred well before the KM3NeT neutrino event, making a direct temporal association unlikely, the unknown redshift of the potential counterpart NVSS J060639–063421 prevents definitive conclusions about cascade contributions. Further constraints on the source’s distance and variability, especially in multiwavelength data, is necessary to pin down the nature of this emission.
Finally, we note that our spectral analysis assumes sources at sufficient redshift ($z\gtrsim 0.1$) for fully developed electromagnetic cascades with characteristic hard spectra ($\Gamma\sim 0.9$), as predicted in [26]. For sources at lower redshift ($z\lesssim 0.1$), the cascade development would be incomplete, producing softer spectral indices that could be consistent with the observed values we measure. However, conventional hadronic accelerators capable of producing >>100 PeV neutrinos at very low redshift ($z\lesssim 0.01$) would need to be extraordinarily luminous. Such nearby powerful sources would be expected to have prominent counterparts across multiple wavelengths, including bright radio, X-ray, and optical emission — which we have not yet observed for our candidate sources.
@@ 4.1 Assumptions in Modeling the Cascade
The simplifying model we adopt here to compute the $\gamma$-ray spectra is limited by a few assumptions. While the cascades become co-linear with the primary emission in the limit of vanishing magnetic fields, large magnetic fields can push the cascades out to several tens of degrees. In modeling the electromagnetic cascades, we assume that the angular extension of the cascade is smaller than $\sim 1^{\circ}$, which is true for time delays that are small enough to observe on $\mathcal{O}(1-10)$ yr time scales. We also assume a simple redshift dependence of the magnetic field, which becomes important for redshifts greater than about $z\sim$1. However, our strongest results are found at smaller redshifts, where this dependence has vanishing effects. Finally, we adopt the EBL model of Saldana-Lopez, et al. [29]. We find that using either the $\pm 1\sigma$ fits to the EBL from this model does not significantly change our results.
@@ 4.2 The True Ultra-High Energy Neutrino Flux
Throughout this paper, we assume that the KM3NeT event was produced by a putative source that not only produced one event (kk=1), but also has a neutrino flux which was expected to produce one event ($\lambda$=1). This approach follows the recent work of Ref. [34], which employed the same methodology to calculate the expected GeV and TeV $\gamma$-ray signals.
@@ 5 Conclusions
[…]